# Beyond the Electric Dipole Approximation: Electric and Magnetic Multipole Contributions Reveal Biaxial Water Structure from SFG Spectra at the Air-Water Interface

### v2

@@ Appendix E Electric Dipole SFG Contribution as a Fingerprint for Interfacial Structure
The second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ plays a central role in SFG theory as it allows to relate the macroscopic electric dipole contribution to the molecular orientation [27, 25, 28, 29]. We can relate $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ to the second-order response function by
where $f^{\alpha}_{i}(z)$ are the laterally averaged local field factors, defined in Equation (1). This is the formalism traditionally used in SFG theory [68, 41, 24, 25, 31, 42]. Note that we do not introduce any locality approximations in the computation of the second-order response $\tilde{s}^{(2)}_{ijk}(z)$, but rather predict $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ by dividing the nonlocal second-order response profile $\tilde{s}^{(2,\mathrm{DD})}_{ijk}\left(z\right)$ by the averaged local field factors $f^{\alpha}_{i}(z)$. We neglect the frequency dependence of $f^{\alpha}_{i}(z)$ by approximating $f^{\alpha}_{i}(z)\approx f^{\mathrm{VIS}}_{i}(z)$. We compare $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ with $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}\left(z\right)$, which is solely based on the anisotropic orientation of the molecules.
@@
The Euler angles $\phi$, $\theta$ and $\psi$ specify the molecule’s orientation and we employ the $z^{\prime}y^{\prime}z^{\prime}$ convention [118]. We define the molecular Eckart frame [110, 111, 119, 120] by aligning the molecular $z^{\prime}$-axis with the permanent dipole vector. The molecular $x^{\prime}$-axis points out of the molecular plane, while the molecular $y^{\prime}$-axis is chosen to be perpendicular to both $x^{\prime}$ and $z^{\prime}$. The molecular Eckart frame is depicted in Figure 6. As shown in SM Section VII [60], the molecular hyperpolarizability $\tilde{\beta}_{ijk}(\theta,\psi)$ is determined by three orientation-dependent functions $q_{10}(\theta)$, $q_{30}(\theta)$ and $q_{32}(\theta,\psi)$. The first and third Legendre polynomials [118]
@@
We obtain $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ by spatial averaging all hyperpolarizabilities according to
@@
where $V$ is the volume, $\rho(z)$ is the molecular density and $q_{lm}(z)$ is the z-dependent average of $q_{lm}(\theta^{n},\psi^{n})$ over all molecules. The frequency dependent coefficients $\tilde{\beta}_{ijk}^{lm}$ are presented in Figure 3 (j). A detailed description of the orientation analysis can be found in SM Section VII [60]. The full information about the orientation distribution is determined by the orientation distribution function defined by

### v3

@@ IV.6 Electric Dipole SFG Contribution as a Fingerprint for Interfacial Structure
The second-order electric dipole susceptibility $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ plays a central role for the interpretation of SFG spectra as it allows to interpret the macroscopic electric dipole contribution in terms of molecular orientation [133, 88, 123, 138]. $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ is related to the electric dipole contribution to the SFG response profile via Equation (1), which is a frequently employed formalism to relate macroscopic and microscopic nonlinear quantities [6, 116, 117, 88, 59]. Note that we do not introduce any locality approximations in the computation of the second-order response $\tilde{s}^{(2)}_{ijk}(z)$, but rather predict $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ by dividing the nonlocal second-order response profile $\tilde{s}^{(2,\mathrm{DD})}_{ijk}\left(z\right)$ by the averaged local field factors $f^{\alpha}_{i}(z)$. We neglect the frequency dependence of $f^{\alpha}_{i}(z)$ by approximating $f^{\alpha}_{i}(z)\approx f^{\mathrm{VIS}}_{i}(z)$ for $\alpha=\mathrm{IR}$ and $\alpha=\mathrm{SFG}$.
We compare $\tilde{\chi}^{(2,\mathrm{DL})}_{ijk}\left(z\right)$ with $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}\left(z\right)$, which is solely based on the anisotropic orientation of the molecules.
@@
The Euler angles $\phi$, $\theta$ and $\psi$ specify the molecule’s orientation and we employ the $z^{\prime}y^{\prime}z^{\prime}$ convention [50]. We define the molecular Eckart frame [36, 57, 3, 103] by aligning the molecular $z^{\prime}$-axis with the permanent dipole vector. The molecular $x^{\prime}$-axis points out of the molecular plane, while the molecular $y^{\prime}$-axis is chosen to be perpendicular to both $x^{\prime}$ and $z^{\prime}$. The molecular Eckart frame is depicted in Figure 6. As shown in SI Section VII, the molecular hyperpolarizability $\tilde{\beta}_{ijk}(\theta,\psi)$ is determined by three orientation-dependent functions $q_{10}(\theta)$, $q_{30}(\theta)$ and $q_{32}(\theta,\psi)$. The first and third Legendre polynomials [50]
@@
We obtain $\tilde{\chi}^{(2,\mathrm{ORI})}_{ijk}(z)$ by averaging all hyperpolarizabilities according to
@@
where $V$ is the volume, $\rho(z)$ is the molecular density and $q_{lm}(z)$ is the z-dependent average of $q_{lm}(\theta^{n},\psi^{n})$ over all molecules. The frequency dependent coefficients $\tilde{\beta}_{ijk}^{lm}$ are presented in Figure 3 (j). A detailed description of the orientation analysis can be found in SI Section VII. The full information about the orientation distribution is determined by the orientation distribution function defined by
@@
V Acknowledgements
VI Data Availability
A data subset supporting the findings of this study is available on Code Ocean under [URL]. Additional data are available from the corresponding author upon reasonable request.
VII Code Availability
Code supporting the findings of this study is available on Code Ocean under [URL].